THe area of the floor in Samantha's rectangular room is 120 square feet. The length of the room is eight feet less than twice the width. What are the dimensions of the floor?
Answer by mananth(16949)

let width be x
length = 2x-8
Artea = L*W
Area = x(2x-8)=120
2x^2-8x=120
/2
x^2-4x=60
x^2-4x-60=0
x^2-10x+6x-60=0
x(x-10)+6(x-10)=0
(x-10)(x+6)=0
x= 10 OR -6
ignore the negative value
width = 10 feet
length = 2x-8
2*10-8
20-8
=12
length = 12 feet